# Conditional computation in neural networks: principles and research trends

### v1

@@ Abstract
Note: This tutorial paper is currently under review at Intelligenza Artificiale (IOS Press). Feedback can be submitted to simone.scardapane@uniroma1.it.
@@ 1 Introduction
Among these design principles, sequentiality has remained a key component. NNs, be they convolutional, recurrent, or transformers, are composed of a stack of differentiable operations, which are activated in sequence for each input to be processed. Stated in another way, their computational graph, i.e., the sequence of primitive operations executed on the underlying hardware, is fixed beforehand when instantiating them. Because NNs have continued to scale in depth and width, this has led to several issues in terms of computational performance and efficiency [22]. Standard techniques to make NNs more efficient only address this partially, by replacing the original network with other static models having less layers (distillation [94], pruning [42]), less precision per parameter, or by approximating each weight matrix (e.g., via low-rank factorization).
@@
The benefits of developing networks that can dynamically adjust their computational graph goes beyond memory or time efficiency. More and more, neural networks are treated in the same way as software [40], and deploying them requires the possibility of quickly debugging their predictions [11], continually fine-tuning them on new data, and transferring parts of their knowledge from one network to the other [3], similarly to standalone software libraries. As we will see, dynamically-activated neural networks provide a principled way to improve both zero-shot transfer and generalization [69] (Sec. 4.2) and explainability of the models (Section 4.3). This aligns with the requirements of many novel applications of NNs, from scientific discovery [92] to AI-native telecommunications [98, 85]. In particular, smart semantic communication networks [85] envisioned in the so-called beyond 6G model necessitate networks that can flexibly adapt to bandwidth and energy constraints, while ensuring transferability and communication through separate neural modules.
@@ 2.2 Discrete sampling: the Gumbel-Softmax trick
Before proceeding, we discuss briefly the implementation of the subsampling layer $\Gamma$. As a prototypical example, we restrict our analysis to subsampling a single element from the set in input. In this case, a very common implementation is the so-called Gumbel-Softmax (GS) trick [35], also known as the concrete distribution [49]. In this section we only provide a high-level overview, and we refer to [34] for a fuller exposition.
Figure 4: Overview of the Gumbel-Softmax trick. We show in orange differentiable operations (the argmax’s gradient being zero almost everywhere), and with dashed arrow the relaxed backward path.
@@
Removing the Gumbel noise $g_{i}$, this corresponds to taking the element with highest score $p_{i}$, while sampling provides a degree of freedom that is helpful in exploring possible alternatives. Since the probabilities $p_{i}$ are implicitly trained via $p(\mathcal{S},x)$, the network can learn to select the element from $\mathcal{S}$ which is most useful for the specific input $x$. In the case of channel selection, for example, this can lead to specializing single channels to specific types of inputs.
@@ 3 Concrete implementations
The previous section has introduced a generic framework for designing networks with dynamic computational graphs. In the literature, these ideas have coalesced around a few, notable set of implementations. In this section we overview three of the most popular ones: early exits (Sec. 3.1), MoE layers (Sec. 3.2), and token sampling (Sec. 3.5) as examples of networks having dynamic width, depth, and input, respectively. In all cases we focus on key strengths and drawbacks of each approach, highlighting how they connect to the framework of Sec. 2.
@@ 3.1 Early exits
Early exit neural networks (EENNs) were introduced based on the idea that not all the inputs to a network require to go through all the layers of the model to be correctly classified [88]. In fact, the accuracy of a network can decrease with respect to the depth on particularly easy samples, a phenomenon known as overthinking [79].
@@
where $\circ$ denotes function composition, $b$ is the last layer of the backbone, and we remove the dependency on the parameters for simplicity. Where to place the auxiliary classifiers is in general an hyper-parameter [81], but here, for clarity, we add a classifier after each block, producing multiple prediction functions:
@@
with the subsampling block over the indexes of the early exits. This can also be given a Bayesian interpretation under the stick-breaking process [68]. While the sampler in (17) is conditioned on the full input, other choices are possible. For example, in multitask learning a separate task embedding can be used to select different exits for different tasks [106], while in the case of transformers only the class token can be used to decide an optimal exit strategy [30]. Being able to train the choice of early exit is fundamental because it allows us to regularize it further for, e.g., better inference-accuracy trade-offs: we revisit this idea in Section 4.1. See [80, 50] for in-depth reviews on EENNs and [80] for a fuller description of differentiable branching.
@@ 3.2 Mixture-of-experts
As stated earlier, the seminal idea of MoEs dates back to [39], where this paradigm is regarded as an ensembling technique with soft gates. The first work that adopts MoEs in a modern deep learning scenario is [83]. Here, the gate becomes sparse, i.e., a fixed top-$k$ operation and the main goal is scaling the number of parameters. Based on these initial results, in the past few years the deep learning community witnessed a steep growth of MoE’s success in conjuction with transformer architectures.
Unlike early exiting, where the routing function typically works at the level of input samples (e.g., images for a CNN), a MoE layer as in (18) can work at the level of individual tokens in a transformer. More specifically, in the fully-connected layers of a transformer block, each token can be assigned to one or more experts. Transformer MoEs have been proposed for NLP [21, 71], for vision [77] and even for multimodal training [57]. Additional variants can be applied to individual heads of an attention block [107] or adapters in a fine-tuning context [105].
@@ 3.4 Soft routing
We discuss here a variant of the MoE layer in (18), known as soft MoE, which has become popular recently to avoid training instabilities in the layer. The key idea is to use the gating function to combine the layer’s inputs [70] or weights [56] instead of the model’s outputs. These variants also provide a fixed compute budget which is decoupled from the total number of parameters, but they are generally easier to implement and to the train.
@@ 3.5 Token dropping and token merging
More formally, let $\mathbf{X}$ be the set of $n$ tokens inside a transformer model stacked horizontally. For the purpose of this section. the subsampling operation $\Gamma$ can be implemented as a matrix multiplication of the input matrix $\mathbf{X}$ and a mask $\mathbf{M}$:
@@ 4.1 Efficiency in training and inference
Conditional computation has gained significant attention for accelerating training and inference in deep learning models [22, 66]. The approaches can vary based on several factors. For example, some research investigated the sparsification of CNNs [90] and of transformers [36, 12], some others focused on some specific aspects of the net, such as the gradients [19], the backpropagation, the activations [13] and the attention layers [89, 14]. For clarity, we focus here on efficiency aspects related to the three families of models seen in Sec. 3, focusing on emerging trends and open challenges.
@@
When training from scratch an EENN, the joint training described in Sec. 3.1 cannot provide faster training since all exits must be trained simultaneously. Soft combinations of the early exits [80] combined with sampling tricks could be useful, but this aspect has been scarcely explored in the literature, possibly due to training instabilities and collapse. Still, EEs can provide indirect benefits to training by accelerating the convergence of deep neural networks [8, 88]. They can also improve the inference efficiency of large-scale pre-trained models [25, 110], making them more discriminative [88], act as a regularization technique [79], and possibly reduce training problems (e.g. vanishing gradient phenomena) [62]. In addition, they can be trained in a layerwise fashion if one starts from a pre-trained network [24].
@@
Compared to EEs, MoEs were investigated mostly to speed-up training by distributing the experts across different GPUs [20], or to accelerate inference by activating a subset of experts for each forward pass. For the former, tiling the model across separate machines requires customized implementations [73]. In addition, to achieve an effective stable throughput, training must avoid collapses of the routing function, and care must be taken to balance the number of tokens which are sent to the different experts. This last point can be achieved with the addition of so-called balancing losses, e.g., batch prioritized routing [77, 15, 20]. Due to these problems, open-source implementations of large MoE models have generally lagged behind proprietary models, with some preliminary steps being taken in this direction [101].
@@
The benefits of applying MoEs can also vary depending on the specific component which is being replaced. In [83], MoEs replace MLP layers to scale-up transformers. A similar methodology, for vision, was presented in [77]. Other approaches proposed to replace attention layers [107], entire blocks [86] of the net, and adapters [105]. By comparison, soft MoEs (Sec. 3.4) can reach similar gains as standard MoE layers while being simpler to train. Finally, the majority of MoEs are trained from scratch, but recently moefication has emerged as an interesting research direction, in which a pre-trained model is converted to a MoE variant by clustering the activations [72, 108]. These variants provide different accuracy-time trade-offs based on the specific routing function. Additional emerging research trends are exploring dynamic variants of routing to provide a flexible inference budget instead [67, 95].
@@
Finally, reducing the number of tokens with token selection techniques is a straightforward strategy to improve the inference time of the network. Most methods in this sense have focused on token dropping [27, 103, 9]: AdaViT [51] proposes a light-weight decision network, attached to the backbone to produce decisions on-the-fly; DynamicViT [75] proposes an attention-masking strategy to block interaction of redundant tokens; MSViT [28] proposes a conditional gating mechanism that selects the token scale for every image region; GTP-ViT [99] introduces a Graph-based Token Propagation (GTP) to propagate less significant tokens’ information to spatially and semantically connected important tokens. During training, token selection is generally achieved by masking the corresponding input elements (Sec. 3.5). Designing token selection mechanisms that can improve training time (for a fixed compute budget) is still an open challenge.
@@
Some works have focused more on specific tasks, such as diffusion [10], long-input sequences [1], or to better understand the general pattern of token dropping [27]. For token merging, PatchMerger [76] proposes a module that reduces the number of tokens by merging them between two consecutive intermediate layers [76]. Recently, a hybrid solution, Token Fusion (ToFU) [42], proposed to put together the benefits of both token pruning and token merging. Token selection can also be combined with the other methods discussed in this paper. As an example, Adaptive Computation Module (ACM) [95] is a technique that combines token dropping with a dynamic width principle, in which each token is allocated a variable width for each layer in the network.
@@ 4.2 Specialization
Studies on quantifying the quality of the resulting specializations were carried out recently in the literature [101]. MoEs showed promising performance due to their nature, and, specifically, hard routing facilitates module specialization [66]. On the other hand, learned routing could lead to sub-optimal results [54, 53] w.r.t to fixed routing, except for specific cases [69]. As another example, [101] showed that many routing decisions tend to ignore context, and they can be fixed during the early stages of training. Generally speaking, understanding the interplay between routing and specialization and the extent to which this specialization correlates with human-understandable semantics and biological plausibility remain open challenges that will require novel benchmarks and metrics [72, 109]. Recent moefications work show that some form of emergent modularity may be found also in pre-trained networks with no specific modularity bias [72, 108].

### v2

@@ Abstract
Please cite the published version as Scardapane, S. et al., 2024. Conditional computation in neural networks: principles and research trends. Intelligenza Artificiale, in press, pp. 1-16. Feedback can be submitted to simone.scardapane@uniroma1.it.
@@ 1 Introduction
Among these design principles, sequentiality has remained a key component. NNs, be they convolutional, recurrent, or transformers, are composed of a stack of differentiable operations, which are activated in sequence for each input to be processed. Stated in another way, their computational graph, i.e., the sequence of primitive operations executed on the underlying hardware, is fixed beforehand when instantiating them. Because NNs have continued to scale in depth and width, this has led to several issues in terms of computational performance and efficiency [22]. Standard techniques to make NNs more efficient only address this partially, by replacing the original network with other static models having fewer layers (distillation [96], pruning [42]), less precision per parameter, or by approximating each weight matrix (e.g., via low-rank factorization).
@@
The benefits of developing networks that can dynamically adjust their computational graph go beyond memory or time efficiency. More and more, neural networks are treated in the same way as software [40], and deploying them requires the possibility of quickly debugging their predictions [11], continually fine-tuning them on new data, and transferring parts of their knowledge from one network to the other [3], similarly to standalone software libraries. As we will see, dynamically-activated neural networks provide a principled way to improve both zero-shot transfer and generalization [71] (Sec. 4.2) and explainability of the models (Section 4.3). This aligns with the requirements of many novel applications of NNs, from scientific discovery [94] to AI-native telecommunications [100, 87]. In particular, smart semantic communication networks [87] envisioned in the so-called beyond 6G model necessitate networks that can flexibly adapt to bandwidth and energy constraints, while ensuring transferability and communication through separate neural modules.
@@ 2.2 Discrete sampling: the Gumbel-Softmax trick
Before proceeding, we discuss briefly the implementation of the subsampling layer $\Gamma$. As a prototypical example, we restrict our analysis to subsampling a single element from the set in input. In this case, a very common implementation is the so-called Gumbel-Softmax (GS) trick [35], also known as the concrete distribution [51]. In this section, we only provide a high-level overview, and we refer to [34] for a fuller exposition.
Figure 4: Overview of the Gumbel-Softmax trick. We show in orange differentiable operations (the argmax’s gradient being zero almost everywhere) and with a dashed arrow the relaxed backward path.
@@
Removing the Gumbel noise $g_{i}$ corresponds to taking the element with the highest score $p_{i}$, while sampling provides a degree of freedom that is helpful in exploring possible alternatives. Since the probabilities $p_{i}$ are implicitly trained via $p(\mathcal{S},x)$, the network can learn to select the element from $\mathcal{S}$ which is most useful for the specific input $x$. In the case of channel selection, for example, this can lead to specializing single channels to specific types of inputs.
@@ 3 Concrete implementations
The previous section has introduced a generic framework for designing networks with dynamic computational graphs. In the literature, these ideas have coalesced around a few, notable set of implementations. In this section we overview three of the most popular ones: early exits (Sec. 3.1), MoE layers (Sec. 3.2), and token sampling (Sec. 3.5) as examples of networks having dynamic depth, width and input, respectively. In all cases we focus on the key strengths and drawbacks of each approach, highlighting how they connect to the framework of Sec. 2.
@@ 3.1 Early exits
Early exit neural networks (EENNs) were introduced based on the idea that not all the inputs to a network are required to go through all the layers of the model to be correctly classified [90]. In fact, the accuracy of a network can decrease with respect to the depth on particularly easy samples, a phenomenon known as overthinking [81].
@@
where $\circ$ denotes function composition, $b$ is the last layer of the backbone, and we remove the dependency on the parameters for simplicity. Where to place the auxiliary classifiers is in general a hyper-parameter [83], but here, for clarity, we add a classifier after each block, producing multiple prediction functions:
@@
with the subsampling block over the indexes of the early exits. This can also be given a Bayesian interpretation under the stick-breaking process [70]. While the sampler in (17) is conditioned on the full input, other choices are possible. For example, in multitask learning a separate task embedding can be used to select different exits for different tasks [108], while in the case of transformers, only the class token can be used to decide an optimal exit strategy [30]. Being able to train the choice of early exit is fundamental because it allows us to regularize it further for, e.g., better inference-accuracy trade-offs: we revisit this idea in Section 4.1. See [82, 52] for in-depth reviews on EENNs and [82] for a fuller description of differentiable branching.
@@ 3.2 Mixture-of-experts
As stated earlier, the seminal idea of MoEs dates back to [39], where this paradigm is regarded as an ensembling technique with soft gates. The first work that adopts MoEs in a modern deep learning scenario is [85]. Here, the gate becomes sparse, i.e., a fixed top-$k$ operation and the main goal is scaling the number of parameters. Based on these initial results, in the past few years, the deep learning community witnessed a steep growth of MoE’s success in conjunction with transformer architectures.
Unlike early exiting, where the routing function typically works at the level of input samples (e.g., images for a CNN), an MoE layer as in (18) can work at the level of individual tokens in a transformer. More specifically, in the fully-connected layers of a transformer block, each token can be assigned to one or more experts. Transformer MoEs have been proposed for NLP [21, 73], for vision [79] and even for multimodal training [59]. Additional variants can be applied to individual heads of an attention block [109] or adapters in a fine-tuning context [107].
@@ 3.4 Soft routing
We discuss here a variant of the MoE layer in (18), known as soft MoE, which has become popular recently to avoid training instabilities in the layer. The key idea is to use the gating function to combine the layer’s inputs [72] or weights [58] instead of the model’s outputs. These variants also provide a fixed compute budget which is decoupled from the total number of parameters, but they are generally easier to implement and train.
@@ 3.5 Token dropping and token merging
More formally, let $\mathbf{X}$ be the set of $n$ tokens inside a transformer model stacked horizontally. For the purpose of this section, the subsampling operation $\Gamma$ can be implemented as a matrix multiplication of the input matrix $\mathbf{X}$ and a mask $\mathbf{M}$:
@@ 4.1 Efficiency in training and inference
Conditional computation has gained significant attention for accelerating training and inference in deep learning models [22, 68]. Often, the time saving comes with a drop in accuracy, making it important to assess the accuracy/computation trade-off. An interesting example is shown in Fig. 5. The approaches can vary based on several factors. For example, some research investigated the sparsification of CNNs [92] and of transformers [36, 12], some others focused on some specific aspects of the net, such as the gradients [19], the backpropagation, the activations [13] and the attention layers [91, 14]. For clarity, we focus here, mainly, on efficiency aspects related to the three families of models seen in Sec. 3, focusing on emerging trends and open challenges. To make the analysis more timely and precise, we have reported all the methods analyzed in Table 1.
@@
When training from scratch an EENN, the joint training described in Sec. 3.1 cannot provide faster training since all exits must be trained simultaneously. Soft combinations of the early exits [82] combined with sampling tricks could be useful, but this aspect has been scarcely explored in the literature, possibly due to training instabilities and collapse. Still, EEs can provide indirect benefits to training by accelerating the convergence of deep neural networks [8, 90]. They can also improve the inference efficiency of large-scale pre-trained models [25, 112], making them more discriminative [90], act as a regularization technique [81], and possibly reduce training problems (e.g. vanishing gradient phenomena) [64]. In addition, they can be trained in a layerwise fashion if one starts from a pre-trained network [24]. In Table 1, we can see that most of these methods require additional parameters, leading to a latency reduction.
@@
Compared to EEs, MoEs were investigated mostly to speed up training by distributing the experts across different GPUs [20], or to accelerate inference by activating a subset of experts for each forward pass. For the former, tiling the model across separate machines requires customized implementations [75]. In addition, to achieve an effective stable throughput, training must avoid collapses of the routing function, and care must be taken to balance the number of tokens that are sent to the different experts. This last point can be achieved with the addition of so-called balancing losses, e.g., batch prioritized routing [79, 15, 20]. Due to these problems, open-source implementations of large MoE models have generally lagged behind proprietary models, with some preliminary steps being taken in this direction [103], as observable also in Table 1.
@@
The benefits of applying MoEs can also vary depending on the specific component that is being replaced. In [85], MoEs replace MLP layers to scale-up transformers. A similar methodology, for vision, was presented in [79]. Other approaches proposed to replace attention layers [109], entire blocks [88] of the net, and adapters [107]. By comparison, soft MoEs (Sec. 3.4) can reach similar gains as standard MoE layers while being simpler to train. Finally, the majority of MoEs are trained from scratch, but recently moefication has emerged as an interesting research direction, in which a pre-trained model is converted to an MoE variant by clustering the activations [74, 110] (see Table 1 to see the approaches based on pre-training). These variants provide different accuracy-time trade-offs based on the specific routing function. Additional emerging research trends are exploring dynamic variants of routing to provide a flexible inference budget instead [69, 97].
@@
Finally, reducing the number of tokens with token selection techniques is a straightforward strategy to improve the inference time of the network. Most methods in this sense have focused on token dropping [27, 105, 9]: A-ViT adopts a token halting approach (shown also in Fig. 5);AdaViT [53] proposes a light-weight decision network, attached to the backbone to produce decisions on-the-fly; DynamicViT [77] proposes an attention-masking strategy to block interaction of redundant tokens; MSViT [28] proposes a conditional gating mechanism that selects the token scale for every image region; GTP-ViT [101] introduces a Graph-based Token Propagation (GTP) to propagate the information of less significant tokens. During training, token selection is generally achieved by masking the corresponding input elements (Sec. 3.5). Designing token selection mechanisms that can improve training time (for a fixed compute budget) is still an open challenge.
@@
Some works have focused more on specific tasks, such as diffusion [10], long-input sequences [1], segmentation [50], or to better understand the general pattern of token dropping [27]. For token merging, PatchMerger [78] proposes a module that reduces the number of tokens by merging them between two consecutive intermediate layers [78]. Recently, a hybrid solution, Token Fusion (ToFu) [42], proposed to put together the benefits of both token pruning and token merging. Token selection can also be combined with the other methods discussed in this paper. As an example, Adaptive Computation Module (ACM) [97] is a technique that combines token dropping with a dynamic width principle, in which each token is allocated a variable width for each layer in the network.
Table 1: Overview of methods to improve efficiency of the models, including baseline algorithms not discussed in the text (e.g., pruning). Parameters: whether the method adds additional parameters. Pre-training: whether the method requires a pre-training phase. Latency/FLOPs: whether the method optimizes one of these two metrics
Method Approach Parameters Pre-training Latency FLOPs
Dynamic Convolutions [92] (2020) Sparsity Yes - Yes Yes
Scaling Transformers [36] (2021) Sparsity - - Yes -
SViTE [12] (2021) Pruning - - Yes Yes
Sparse Momentum [19] (2019) Pruning - - Yes Yes
SparseViT [13] (2023) Pruning Yes - Yes Yes
Sparse Transformers [14] (2019) Factorization Yes - - Yes
Differentiable branching [82] (2020) Early Exits Yes - Yes Yes
Adaptive Early Exits [8] (2017) Early Exits Yes - Yes -
L2W-DEN [25] (2022) Early Exits Yes - Yes -
PABEE [112] (2020) Early Exits Yes - Yes -
Branchynet [90] (2016) Early Exits Yes - Yes -
AEP [81] (2023) Early Exits Yes - Yes Yes
BoF [64] (2020) Early Exits Yes - - Yes
DeepSpeed-MoE[75] (2022) MoE Yes Yes Yes Yes
Sparsely-Gated MoE [85] (2017) MoE Yes - - Yes
V-MoE [79] (2021) MoE Yes - - Yes
MoA [109] (2022) MoE Yes - - Yes
SUT [88] (2023) MoE Yes - - Yes
MoV/MoLoRA [107] (2023) MoE Yes Yes - Yes
EMoE [74] (2023) MoE - Yes - Yes
GMoE [46] (2022) MoE Yes Yes - Yes
MoEfication [110] MoE Yes Yes - Yes
SADMoE [69] (2023) MoE Yes Yes - -
A-ViT [105] (2022) Token Sampling - - Yes Yes
ToMe [9, 10] (2022) Token Sampling Yes - Yes Yes
AdaViT [53] (2022) Token Sampling Yes - - Yes
DynamicViT [77] (2021) Token Sampling Yes - Yes Yes
MSViT [28] (2023) Token Sampling Yes - Yes Yes
GTP-ViT [101] (2024) Token Sampling Yes - Yes Yes
Colt5 [1] (2023) MoE Yes Yes Yes Yes
PatchMerger [78] (2022) Token Sampling Yes Yes Yes Yes
ToFu [42] (2024) Token Sampling - Yes Yes Yes
ACM [97] (2023) MoE/Token Sampling Yes Yes - Yes
Figure 5: Example of accuracy-flops trade-off for inference with A-ViT [105]. Specifically, the architecture is a DeiT, trained on Imagenette.
@@ 4.2 Specialization
Studies on quantifying the quality of the resulting specializations were carried out recently in the literature [103]. MoEs showed promising performance due to their nature, and, specifically, hard routing facilitates module specialization [68]. On the other hand, learned routing could lead to sub-optimal results [56, 55] w.r.t to fixed routing, except for specific cases [71]. As another example, [103] showed that many routing decisions tend to ignore context, and they can be fixed during the early stages of training. Generally speaking, understanding the interplay between routing and specialization and the extent to which this specialization correlates with human-understandable semantics and biological plausibility remain open challenges that will require novel benchmarks and metrics [74, 111]. Recent moefications works show that some form of emergent modularity may also be found in pre-trained networks with no specific modularity bias [74, 110].
@@ 5 Conclusions and future trends
Acknowledgments
S. Scardapane is partly funded by Sapienza grants RM1221816BD028D6 (DESMOS) and RG123188B3EF6A80 (CENTS). P. Minervini is partially funded by ELIAI (The Edinburgh Laboratory for Integrated Artificial Intelligence), EPSRC (grant no. EP/W002876/1), an industry grant from Cisco, and a donation from Accenture LLP.